# Reevaluating Policy Gradient Methods for Imperfect-Information Games

### v1

@@ Abstract
In the past decade, motivated by the putative failure of naive self-play deep reinforcement learning (DRL) in adversarial imperfect-information games, researchers have developed numerous DRL algorithms based on fictitious play (FP), double oracle (DO), and counterfactual regret minimization (CFR). In light of recent results of the magnetic mirror descent algorithm, we hypothesize that simpler generic policy gradient methods like PPO are competitive with or superior to these FP, DO, and CFR-based DRL approaches. To facilitate the resolution of this hypothesis, we implement and release the first broadly accessible exact exploitability computations for four large games. Using these games, we conduct the largest-ever exploitability comparison of DRL algorithms for imperfect-information games. Over 5600 training runs, we find that FP, DO, and CFR-based approaches fail to outperform generic policy gradient methods.
Machine Learning, ICML
@@ 1 Introduction
A natural approach to model-free deep reinforcement learning (DRL) in such games is to deploy a single-agent algorithm in self-play. However, because imperfect information induces cyclical best response dynamics, such an approach can fail catastrophically, yielding policies that are maximally exploitable. So as to avoid this outcome, most existing literature focuses on adapting pre-existing tabular algorithms with established equilibrium convergence guarantees, such as fictitious play (FP) (fp51; robinson_fp51), double oracle (DO) (do03), and counterfactual regret minimization (CFR) (cfr07), to DRL.
Unfortunately, constructing effective DRL approaches from these tabular algorithms has proven challenging. FP and DO require expensive best response computations at each iteration (i.e., each iteration requires solving an entire reinforcement learning problem) and can exhibit slow convergence as a function of iteration number (do_lower24). While CFR converges more quickly, its immediate adaptation to model-free reinforcement learning requires importance weighting over trajectories (mccfr09), resulting in high variance feedback to which function approximation is poorly suited. Moreover, none of FP, DO and CFR enjoy last-iterate convergence, creating a layer of indirection to extract the desired policy.
Recently, mmd23 demonstrated the promise of an alternative algorithm, a policy gradient (PG) method called magnetic mirror descent (MMD). They showed that MMD achieves competitive performance with CFR in tabular settings, while retaining the deep learning compatibility native to reinforcement learning algorithms.
In this work, we expound the relationship between MMD and generic deep PG methods such as PPO (ppo17). Like MMD, the improvement steps of these other generic PG methods 1) maximize expected value, 2) regularize the policy, and 3) control the size of the update. These parallels lead us to the question: Given that MMD performs well as a DRL algorithm for IIGs, shouldn’t these other generic PG methods perform well, too?
We believe that the answer to this question is yes.
To further this idea, we put forth the following hypothesis:
With proper tuning, generic PG methods are
competitive with or superior to FP, DO, and CFR-based DRL approaches in IIGs.
@@
However, assessing the above hypothesis is not straightforward. Doing so demands exact exploitability computations for games large enough to be representative of settings to which DRL is relevant, but small enough that these computations are tractable in reasonable time. Unfortunately, due to the complexity of implementing these computations for large games, few are publicly available, even in OpenSpiel (openspiel19)—the standard library for DRL in IIGs. This issue has plagued the field, forcing researchers to resort to tenuous metrics and making progress difficult to measure.
We address this issue by implementing exact exploitability computation for four large games—3x3 Dark Hex, Phantom Tic-Tac-Toe, 3x3 Abrupt Dark Hex, and Abrupt Phantom Tic-Tac-Toe—ourselves. These games have 10s of billions of nodes in their game trees—over one million times more than many games for which exploitability is typically reported, such as Leduc Hold’em (leduc05).
Armed with these benchmarks, we undertake the largest-ever exploitability comparison of DRL algorithms for IIGs, spanning 5600 total runs across 350 hyperparameter configurations, requiring over 278,000 CPU hours, and including NFSP (nfsp16), PSRO (psro17), ESCHER (escher23), R-NaD (deepnash22), MMD (mmd23), PPO (ppo17) and PPG (ppg20). Over these runs, all of which are included in the paper, NFSP, PSRO, ESCHER, and R-NaD fail to outperform the generic PG methods (MMD, PPO, PPG).
We release our code for computing exact exploitability and head-to-head values, which require under 2 minutes on commodity hardware, with OpenSpiel-compatible Python bindings, as well as efficient implementations of tabular solvers and best responses. We encourage the community to use these tools both to test our hypothesis and as benchmarks for future research.
@@ 2 Preliminaries
To set context, we give a formalization of IIGs and exploitability, and an overview of model-free DRL approaches to IIGs.
@@ 2.1 Game Formalism
IIGs are often formalized as perfect-recall extensive-form games (EFGs) (kuhn2003lectures). Our formalism is equivalent to perfect-recall EFGs, but possesses greater superficial similarity to traditional reinforcement learning notation. We notate a game as a tuple
@@
  • $t_{\text{max}}$ is the time step after which the game terminates.
The players interact with the game using policies $\pi_{i}\colon\mathbb{I}\to\Delta(\mathbb{A})$ for $i\in\{0,1\}$ mapping information states to distributions over actions.
The objective of player 1 and loss of player 2 is the expected return:
$\mathcal{J}(\pi)=\mathbb{E}_{\pi}\left[G\right]=\mathbb{E}_{\pi}\left[\sum_{t=1}^{t_{\text{max}}}\mathcal{R}(S^{t},A^{t})\right].$
However, directly optimizing the expected return is only possible given a particular opponent or distribution of opponents. In practice, identifying this distribution may be prohibitively expensive or infeasible—such a distribution may not even exist.
@@
$\text{expl}(\pi)=\frac{\max_{\pi_{0}^{\prime}}\mathcal{J}(\pi_{0}^{\prime},\pi_{1})-\min_{\pi_{1}^{\prime}}\mathcal{J}(\pi_{0},\pi_{1}^{\prime})}{2}.$
Exploitability is the expected return of a worst-case opponent, called a best response, assuming it plays half of the time as player 1 and half of the time as player 2. Joint policies with exploitability zero are Nash equilibria.
@@ 2.2 Algorithmic Approaches
Existing model-free DRL approaches to two-player zero-sum IIGs are largely based on one of naive self-play, best responses, CFR, and regularization.
@@ 2.2.1 Naive self-play
Applying model-free DRL algorithms to two-player zero-sum IIGs in self-play is appealing for its simplicity; however, it is easy to show that doing so naively can lead to catastrophic results. Some value-based algorithms cannot even express non-deterministic policies. This makes it impossible for them to achieve low-exploitability in many games (e.g., in rock-paper-scissors, deterministic policies are maximally exploitable). While PG methods can at least express non-deterministic policies, their learning dynamics generally cycle, diverge or exhibit chaotic behavior, rather than converge to Nash equilibria (vortices19).
@@ 2.2.2 Best Responses
A next-simplest alternative is to leverage tabular game-theoretic algorithms that rely on best-response computations. These game-theoretic algorithms support a plug-and-play workflow with model-free DRL, by substituting the exact best response for an approximate best response computed via DRL. And, unlike naive self-play, they come with convergence guarantees.
Fictitious play (FP) (fp51; robinson_fp51) is one such game-theoretic algorithm. At each iteration $t$, FP computes a best response $\pi_{t}$ to the average of the previous iterations $\bar{\pi}_{t-1}=\text{avg}(\pi_{1},\dots,\pi_{t-1})$. This average is over the sequence form of the policy, meaning that it is equivalent to a policy that uniformly samples $T\sim\text{uniform}(\{1,\dots,t-1\})$ at the start of each game and then plays according to $\pi_{T}$ until the game’s conclusion. The exploitability of the average $\bar{\pi}_{t}$ of a sequence of policies generated by FP converges to zero as $t$ grows large.
Double oracle (DO) (do03) is another such game-theoretic algorithm. Like FP, DO works by iteratively computing best responses. However, rather than computing best responses to the average of the previous iterates, DO computes best responses $\pi_{t}$ to Nash equilibria $\sigma_{t}^{\ast}$ of a so-called metagame. This metagame is a one-shot game in which players select policies $\pi^{(1)},\pi^{(2)}$ from among the previous best responses $\{\pi_{1},\dots,\pi_{t-1}\}$ and receive as payoff the expected value $\mathcal{J}(\pi^{(1)},\pi^{(2)})$ of playing these policies against one another. A policy of this metagame is a mixture $\sigma_{t}$ over previous best responses $\{\pi_{1},\dots,\pi_{t-1}\}$ that is enacted by sampling $\pi_{T}\sim\sigma_{t}$ once at the start of each game and then playing according to $\pi_{T}$ until the game’s conclusion. A policy of this metagame $\sigma_{t}$ is unexploitable if none of the previous best responses $\{\pi_{1},\dots,\pi_{t-1}\}$ exploits it. The exploitability of a sequence of metagame Nash equilibria is guaranteed to converge to zero as $t$ grows large.
Unfortunately, both FP and DO can suffer from slow convergence. This is true not only in practice, but also according to the exponential best-known complexity for FP and exponential lower bound for DO (do_lower24). Requiring a large number of iterations is especially concerning in the context of DRL because each iteration requires solving an entire reinforcement learning problem. Despite numerous works on DRL variants of both FP (nfsp16; nfsp-em19; alphastar19; nfsp-adapt23; nfsp-eve23; nfsp-plt23) and DO (psro17; psro-pipe20; xdo21; iterative-psro21; anytimepsro22; odo22; epsro22; psro-nontrans23; rmdo23; sppsro24; sapsro24; fusion-psro24), it is not clear that this issue has been addressed.
@@ 2.2.3 Counterfactual Regret Minimization
Counterfactual regret minimization (CFR) (cfr07), widely regarded as the gold standard for tabularly solving IIGs, generally converges faster than FP or DO. Instead of relying on best responses, CFR is grounded in the principle of regret minimization. A key result of regret minimization is that, in two-player zero-sum games, the average strategy of two regret minimizing learners converges to a Nash equilibrium. CFR applies this principle by independently minimizing regret at each information state. By appropriately weighting feedback, CFR ensures regret minimization across the entire game, thus guaranteeing convergence to a Nash equilibrium.
Because CFR is more involved than FP and DO, it does not admit the same kind of plug-and-play workflow with single-agent model-free DRL. Instead, new methods (dream20; armac20; escher23) have been designed specifically around CFR. Unfortunately, these new methods come with new weaknesses. The most straightforward method (dream20), for instance, relies on importance sampling over trajectories, which leads to high variance in feedback, especially in games with many steps. This makes it challenging for function approximators to learn. Although escher23 resolve this issue, their resolution necessitates a uniform behavioral policy, making it difficult to achieve sufficient coverage of relevant trajectories.
@@ 2.2.4 Regularization
Motivated by the limitations of FP, DO, and CFR-based methods, fforel21 proposed an alternative approach centered around regularization. Unlike previous methods, their approach is designed to converge in the last iterate, aligning it more closely with typical DRL techniques. fforel21’s work has been influential, serving as inspiration for several subsequent studies (fu2022actorcritic; deepnash22; mmd23; ftrl-ent23; reg-enough24). Among them, mmd23 introduced a policy gradient algorithm called magnet mirror descent (MMD), characterized by the following update rule:
@@
where $q(A)$ is the value of action $A$, $\alpha$ is a regularization temperature, $\eta$ is a stepsize, $\rho$ is a magnet policy, and KL is KL divergence. In a surprising result, mmd23 demonstrated that MMD, implemented as a standard tabular self-play reinforcement learning algorithm, achieves performance competitive with that of CFR in IIGs. mmd23 also give empirical evidence that MMD is performant as a DRL algorithm.
@@ 3 The Policy Gradient Hypothesis
  • $-\alpha\text{KL}(\pi,\rho)$ regularizes the updated policy (if $\rho$ is uniform, this functions similarly to an entropy bonus).
@@
Far from being unique to MMD, these functions have long been present in deep PG methods, such as TRPO (trpo15), PPO (ppo17), and PPG (ppg20). Indeed, as with MMD, the update rules of these algorithms include terms that maximize expected value and encourage entropy; they also possess mechanisms that, by various means, control the size of the update.
Yet, despite this shared ethos, generic PG methods—typically PPO—appear in literature only as underperforming baselines. This raises the question: Why has MMD performed well, while these generic PG methods have not?
There are multiple answers worth considering. One possibility is that the differences between MMD and other generic PG methods are material. Unlike most PG methods, which use some combination of gradient clipping and forward KL divergence, MMD employs reverse KL divergence to regulate update size. In tabular settings, where competitive performance requires high precision, this distinction is manifestly important.
However, in the context of deep learning, it is less obvious that this variation matters. Empirical research has raised questions about the extent to which these different approaches to controlling update size vary in performance (Engstrom2020Implementation; revisiting20; forwardreverse22), and suggested that none of them actually succeed in controlling update size (Ilyas2020A).
@@
But there are also several domain-specific factors that make this second possibility particularly plausible:
@@
  • Works that report generic PG method baselines often rely on head-to-head results. The intransitive nature of these results increases the complexity of tuning, and may thereby have hampered it.
  • The hyperparameter regime for which mmd23 showed MMD to be effective involves more entropy regularization than is typically used for PG methods in single-agent problems. Thus, it is possible effective hyperparameter configurations were overlooked due to their regime being too far removed from those of single-agent settings.
@@
The confirmation of the policy gradient hypothesis would have ramifications on both research and practice. On the research side, it would vindicate generic PG methods, which have generally either been dismissed as unsound or relegated to the role of sacrificial baseline. Practically, it would lead to simpler implementations and potentially significant performance improvements in applications where FP, DO, and CFR-based DRL algorithms are currently employed. These applications include autonomous vehicles (nfsp_av18), network security (psro_scale_free20; nfsp_network21), robot confrontation (psro_robot22), eavesdropping (nfsp_eavesdrop22), radar (anti-)jamming (nfsp_radar22; cfr_jamming22; nfsp_radar23; nfsp_antijamming24), intrusion response (nfsp_security23), aerial combat (nfsp_dogfight23; nfsp_dogfight24), pursuit-evasion games (psro_pursuit23; psro_pursuit24), racing (zheng2024racing), language model alignment (psro_lm24), and language model red teaming (psro_red_team24).

### v2

@@ Abstract
In the past decade, motivated by the putative failure of naive self-play deep reinforcement learning (DRL) in adversarial imperfect-information games, researchers have developed numerous DRL algorithms based on fictitious play (FP), double oracle (DO), and counterfactual regret minimization (CFR). In light of recent results of the magnetic mirror descent algorithm, we hypothesize that simpler generic policy gradient methods like PPO are competitive with or superior to these FP-, DO-, and CFR-based DRL approaches. To facilitate the resolution of this hypothesis, we implement and release the first broadly accessible exact exploitability computations for four large games. Using these games, we conduct the largest-ever exploitability comparison of DRL algorithms for imperfect-information games. Over 5600 training runs, we find that FP-, DO-, and CFR-based approaches fail to outperform generic policy gradient methods.
Code available at https://github.com/nathanlct/IIG-RL-Benchmark and https://github.com/gabrfarina/exp-a-spiel.
@@ 1 Introduction
A natural approach to model-free deep reinforcement learning (DRL) in such games is to deploy a single-agent algorithm in self-play. However, because imperfect information induces cyclical best response dynamics, such an approach can fail catastrophically, yielding policies that are maximally exploitable. So as to avoid this outcome, most existing literature focuses on adapting pre-existing tabular algorithms with established equilibrium convergence guarantees, such as fictitious play (FP) (Brown, 1951; Robinson, 1951), double oracle (DO) (McMahan et al., 2003), and counterfactual regret minimization (CFR) (Zinkevich et al., 2007), to DRL.
Unfortunately, constructing effective DRL approaches from these tabular algorithms has proven challenging. FP and DO require an expensive best response computation at each iteration (i.e., each iteration requires solving an entire reinforcement learning problem) and can exhibit slow convergence as a function of iteration number (Zhang and Sandholm, 2024). While CFR converges more quickly, its immediate adaptation to model-free reinforcement learning requires importance weighting over trajectories (Lanctot et al., 2009), resulting in high variance feedback to which function approximation is poorly suited. Moreover, none of FP, DO and CFR enjoy last-iterate convergence, necessitating a layer of indirection to extract the desired policy.
Recently, Sokota et al. (2023) demonstrated the promise of an alternative algorithm, a policy gradient (PG) method called magnetic mirror descent (MMD). They showed that MMD achieves competitive performance with CFR in tabular settings, while retaining the deep learning compatibility native to reinforcement learning algorithms.
In this work, we expound the relationship between MMD and other generic deep PG methods such as PPO (Schulman et al., 2017). Like MMD, the improvement steps of these other generic PG methods 1) maximize expected value, 2) regularize the policy, and 3) control the size of the update. These parallels lead us to the question: Given that MMD performs well as a DRL algorithm for IIGs, shouldn’t these other generic PG methods perform well, too?
We believe that the answer to this question is yes. To further this idea, we put forth the following hypothesis:
With proper tuning, generic PG methods are competitive with or
superior to FP-, DO-, and CFR-based DRL approaches in IIGs.
@@
However, assessing the above hypothesis is not straightforward. Doing so demands exact exploitability computations for games large enough to be representative of settings to which DRL is relevant, but small enough that these computations are tractable in reasonable time. Unfortunately, due to the complexity of implementing these computations for large games, few are publicly available, even in OpenSpiel (Lanctot et al., 2019)—the standard library for DRL in IIGs. This issue has plagued the field, forcing researchers to resort to tenuous metrics and making progress difficult to measure.
We address this issue by implementing exact exploitability computation for four large games: 3x3 Dark Hex, Phantom Tic-Tac-Toe, 3x3 Abrupt Dark Hex, and Abrupt Phantom Tic-Tac-Toe. These games have 10s of billions of nodes in their game trees—over one million times more than many games for which exploitability is typically reported, such as Leduc Hold’em (Southey et al., 2005).
Using these benchmarks, we undertake the largest-ever exploitability comparison of DRL algorithms for IIGs, spanning 5600 total runs across 350 hyperparameter configurations, requiring over 278,000 CPU hours, and including NFSP (Heinrich and Silver, 2016), PSRO (Lanctot et al., 2017), ESCHER (McAleer et al., 2023), R-NaD (Perolat et al., 2022), MMD (Sokota et al., 2023), PPO (Schulman et al., 2017) and PPG (Cobbe et al., 2021). Over these runs, all of which are included in the paper, NFSP, PSRO, ESCHER, and R-NaD fail to outperform the generic PG methods (MMD, PPO, PPG).
We release our implementations for computing exact exploitability and head-to-head values, which require under 2 minutes on commodity hardware, with OpenSpiel-compatible Python bindings, as well as efficient implementations of tabular solvers and best responses. We encourage the community to use these tools both to test our hypothesis and as benchmarks for future research.
@@ 2 Preliminaries
To start, we set context by giving a formalization of IIGs and exploitability, and an overview of model-free DRL approaches to IIGs.
@@ 2.1 Game Formalism
IIGs are often formalized as perfect-recall extensive-form games (EFGs) (Kuhn, 2003). Our formalism is equivalent to perfect-recall EFGs, but possesses greater superficial similarity to traditional reinforcement learning notation. We notate a game as a tuple
@@
  • $t_{\text{max}}$ is the time step at which the game terminates.
The players interact with the game using policies $\pi_{i}\colon\mathbb{I}\to\Delta(\mathbb{A})$ for $i\in\{1,2\}$ mapping information states to distributions over actions.
The objective of player 1 and loss of player 2 is the expected return
$\mathcal{J}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=1}^{t_{\text{max}}}\mathcal{R}(S^{t},A^{t})\right].$
However, directly optimizing the expected return is only possible given a particular opponent or distribution of opponents. In practice, identifying this distribution may be prohibitively expensive or altogether infeasible—such a distribution may not even exist.
@@
$\text{expl}(\pi)=\frac{\max_{\pi_{1}^{\prime}}\mathcal{J}(\pi_{1}^{\prime},\pi_{2})-\min_{\pi_{2}^{\prime}}\mathcal{J}(\pi_{1},\pi_{2}^{\prime})}{2}.$
Exploitability is the expected return of a worst-case opponent, called a best response, playing half of the time as player 1 and half as player 2. Joint policies with exploitability zero are Nash equilibria.
@@ 2.2 Algorithmic Approaches
Existing model-free DRL approaches to producing low exploitability policies for two-player zero-sum IIGs are largely based on one of naive self-play, best responses, CFR, and regularization.
@@ 2.2.1 Naive self-play
Of these approaches, naive self-play is appealing for its simplicity; however, it is easy to show that it can produce catastrophic results. Some value-based algorithms cannot even express non-deterministic policies. This makes it impossible for them to achieve low-exploitability in many games (e.g., in rock-paper-scissors, deterministic policies are maximally exploitable). While PG methods can at least express non-deterministic policies, their learning dynamics generally cycle, diverge or exhibit chaotic behavior, rather than converge to Nash equilibria (Cheung and Piliouras, 2019).
@@ 2.2.2 Best Responses
An alternative is to leverage tabular game-theoretic algorithms that rely on best-response computations. These game-theoretic algorithms support a plug and play workflow with model-free DRL, by substituting the exact best response for an approximate best response computed via DRL. And, unlike naive self-play, they come with convergence guarantees.
Fictitious play (FP) (Brown, 1951; Robinson, 1951) is one such game-theoretic algorithm. At each iteration $t$, FP computes a best response $\pi_{t}$ to the average of the previous iterations $\bar{\pi}_{t-1}=\text{avg}(\pi_{1},\dots,\pi_{t-1})$. This average is over the sequence form of the policy, meaning that it is equivalent to a policy that uniformly samples $T\sim\text{uniform}(\{1,\dots,t-1\})$ at the start of each game and then plays according to $\pi_{T}$ until the conclusion of the game. The exploitability of the average $\bar{\pi}_{t}$ of a sequence of policies generated by FP converges to zero as $t$ grows large.
Double oracle (DO) (McMahan et al., 2003) is another such game-theoretic algorithm. Like FP, DO works by iteratively computing best responses. However, rather than computing best responses to the average of the previous iterates, DO computes best responses $\pi_{t}$ to Nash equilibria $\sigma_{t}^{\ast}$ of a so-called metagame. This metagame is a one-shot game in which players select policies $\pi^{(1)},\pi^{(2)}$ from among the previous best responses $\{\pi_{1},\dots,\pi_{t-1}\}$ and receive as payoff the expected value $\mathcal{J}(\pi^{(1)},\pi^{(2)})$ of playing these policies against one another. A policy of this metagame is a mixture $\sigma_{t}$ over previous best responses $\{\pi_{1},\dots,\pi_{t-1}\}$ that is enacted by sampling $\pi_{T}\sim\sigma_{t}$ once at the start of each game, then playing according to $\pi_{T}$ until the conclusion of the game. A policy of this metagame $\sigma_{t}$ is unexploitable if none of the previous best responses $\{\pi_{1},\dots,\pi_{t-1}\}$ exploits it. The exploitability of a sequence of metagame Nash equilibria is guaranteed to converge to zero as $t$ grows large.
Unfortunately, both FP and DO can suffer from slow convergence. This is suggested not only by practice, but also according to the exponential best-known complexity for FP and exponential lower bound for DO (Zhang and Sandholm, 2024). Requiring a large number of iterations is especially concerning in the context of DRL because each iteration requires solving an entire reinforcement learning problem. Despite numerous works on DRL variants of both FP (Heinrich and Silver, 2016; Qin et al., 2019; Vinyals et al., 2019; Goldstein and Brown, 2023; Hu et al., 2023; Li et al., 2023a) and DO (Lanctot et al., 2017; McAleer et al., 2020, 2021; Smith et al., 2021; McAleer et al., 2022; Dinh et al., 2022; Zhou et al., 2022; Yao et al., 2023; Tang et al., 2023; McAleer et al., 2024; Li et al., 2024a; Lian, 2024), it is not clear that this issue has been addressed.
@@ 2.2.3 Counterfactual Regret Minimization
Counterfactual regret minimization (CFR) (Zinkevich et al., 2007), widely regarded as the gold standard for tabularly solving IIGs, generally converges faster than FP or DO. Instead of relying on best responses, CFR is grounded in the principle of regret minimization. A key result of regret minimization is that, in two-player zero-sum games, the average strategy of two regret minimizing learners converges to a Nash equilibrium. CFR applies this principle by independently minimizing regret at each information state. By appropriately weighting feedback, CFR ensures regret minimization across the entire game, thus guaranteeing convergence to a Nash equilibrium.
Because CFR is more involved than FP and DO, it does not admit the same kind of plug and play workflow with single-agent model-free DRL. Instead, new methods (Steinberger et al., 2020; Gruslys et al., 2020; McAleer et al., 2023) have been designed specifically around CFR. Unfortunately, these new methods come with new weaknesses. The most straightforward method (Steinberger et al., 2020), for instance, relies on importance sampling over trajectories, which leads to high variance in feedback, especially in games with many steps. This makes it challenging for function approximators to learn. Although McAleer et al. (2023) resolve this issue, their resolution necessitates a uniform behavioral policy, making it difficult to achieve sufficient coverage of relevant trajectories.
@@ 2.2.4 Regularization
Motivated by the limitations of FP-, DO-, and CFR-based methods, Perolat et al. (2021) proposed an alternative approach centered around regularization. Unlike previous methods, their approach is designed to converge in the last iterate, aligning it more closely with typical DRL techniques. Perolat et al. (2021)’s work has been influential, serving as inspiration for several subsequent studies (Fu et al., 2022; Perolat et al., 2022; Sokota et al., 2023; Meng et al., 2023; Liu et al., 2024). Among them, Sokota et al. (2023) introduced a policy gradient algorithm called magnet mirror descent (MMD), characterized by the following update rule:
@@
where $q(A)$ is the value of action $A$, $\alpha$ is a regularization temperature, $\eta$ is a stepsize, $\rho$ is a magnet policy, and KL is KL divergence. In a surprising result, Sokota et al. (2023) demonstrated that MMD, implemented as a standard tabular self-play reinforcement learning algorithm, achieves performance competitive with that of CFR in IIGs. Sokota et al. (2023) also gave empirical evidence that MMD is performant as a DRL algorithm.
@@ 3 The Policy Gradient Hypothesis
  • $-\alpha\text{KL}(\pi,\rho)$ regularizes the updated policy (for uniform $\rho$, this acts similarly to an entropy bonus).
@@
Far from being unique to MMD, these functions have long been present in deep PG methods, such as TRPO (Schulman et al., 2015), PPO (Schulman et al., 2017), and PPG (Cobbe et al., 2021). As with MMD, the update rules of these algorithms include terms that maximize expected value and encourage entropy; they also possess mechanisms that, by various means, control the update size.
Yet, despite this shared ethos, other generic PG methods—typically PPO—appear in literature only as underperforming baselines. This raises the question: Why has MMD performed well, while these other generic PG methods have not?
There are multiple answers worth considering. One possibility is that the differences between MMD and other generic PG methods are material. Unlike most PG methods, which use some combination of gradient clipping and forward KL divergence, MMD employs reverse KL divergence to regulate update size. In tabular settings, where competitive performance requires high precision, this distinction is manifestly important. However, in the context of deep learning, it is less clear that this variation matters (Engstrom et al., 2020; Hsu et al., 2020; Chan et al., 2022).
@@
But there are also domain-specific factors that make this second possibility particularly plausible:
@@
  • Works that report generic PG method baselines often rely on head-to-head results. The intransitive nature of these results increases the complexity of tuning and may thereby have hampered it.
  • The hyperparameter regime for which Sokota et al. (2023) showed MMD to be effective involves more entropy regularization than is typically used for PG methods in single-agent problems. Thus, it is possible effective hyperparameter configurations were overlooked due to their regime being too far removed from those of single-agent settings.
@@
The confirmation of the policy gradient hypothesis would have ramifications on both research and practice. On the research side, it would vindicate generic PG methods, which have generally either been dismissed as unsound or relegated to the role of sacrificial baseline. Practically, it would lead to simpler implementations and potentially significant performance improvements in applications where FP-, DO-, and CFR-based DRL algorithms are currently employed. These applications include autonomous vehicles (Ma et al., 2018), network security (Lu et al., 2020; Xue et al., 2021), robot confrontation (Hu et al., 2022), eavesdropping (Guo et al., 2022), radar (anti-)jamming (Li et al., 2022a, b; Geng et al., 2023, 2024), intrusion response (Hammar and Stadler, 2023), aerial combat (He et al., 2023; Wang et al., 2024), pursuit-evasion games (Li et al., 2023b, 2024b), racing (Zheng et al., 2024), language model alignment (Gemp et al., 2024) and red teaming (Ma et al., 2024).